# Bounds on the photon mass via the Shapiro effect in the solar system

### v1

@@ III.3 Bounds from the Cassini-Huygens mission
The best measurement of the Shapiro effect to date was performed using the Cassini spacecraft on its way to Saturn. During its superior conjunction in mid 2002, the spacecraft was at a distance $r_{r}=8.43$ AU and the electromagnetic signal reached its closest approach to the Sun at $r_{0}=1.6\,R_{\odot}$. The radio link was setup so that essentially two carrier signals were sent to and from the spacecraft. The uplink had $f_{u}=7.2$ GHz in the X-band and $f_{u}=34$ GHz in the Ka-band, whereas the downlink had $f_{d}=8.4$ GHz in the X-band and $f_{d}=32$ GHz in the Ka-band. The possibility of using carriers with different frequencies (in particular the Ka-band) allowed the almost complete elimination of coronal effects. Doppler-tracking of the signals via the fractional frequency shifts couuld be used to constrain $\gamma$ to unprecedented accuracy: $|\gamma-1|=2.3\times 10^{-5}$ Cassini_1 ; Cassini_2 .
To simplify the analysis, we will consider the X- and Ka-bands separately. This is also meaningful, as the finite photon mass induces dispersive effects. Since the up- and downlink frequencies are not the same, we are not allowed to use eq. (21). Instead, we must return to eq. (13) and treat the in- and outbound paths differently, assigning to each its frequency. Doing that we find
$\Delta\tau_{\rm delay}$ $=$ $\frac{4GM_{\odot}}{c^{3}}\left[\left(2+\alpha_{\rm u}^{2}+\alpha_{\rm d}^{2}\right)+\log\left(\frac{4r_{e}r_{c}}{r_{0}^{2}}\right)\right]$ (24)
-- $\frac{3GM_{\odot}}{c^{3}}\left(1+\frac{r_{c}}{r_{e}}\right)\,,$
where $\alpha_{\rm u,d}$ indicate the parameter (20) evaluated at frequency $f_{\rm u,d}$. This expression may now be compared with eq. (18), giving us the same sensitivity as in eq. (21), but with the substitution
$f\rightarrow\frac{\sqrt{2}f_{u}f_{d}}{\sqrt{f_{u}^{2}+f_{d}^{2}}}$ (25)
with $f_{u,d}$ in GHz. Using the values quoted above with eqs. (25) and (21), we finally obtain
@@
[…]
The bound (26a) represents a 2.6-fold improvement over the bound from the Viking 1 and 2 data, cf. eq. (23).
III.4 Future perspectives
In the previous section we explored measurements of the gravitational time delay performed by past missions; the last one, Cassini-Huygens, was in the years 2000. Fortunately, there are a few promising proposals, either already underway or planned, that aim at significantly improving the current limits on $\gamma$, among other less well constrained PPN parameters Turyshev . A few examples are the Mercury Orbiter Radioscience Experiment (MORE) MORE ; MORE_1 and the Gravitational Time Delay Mission (GTDM) GTDM which, if all technical requirements are properly met, will be able to measure $|\gamma-1|$ at much higher accuracies than attained by the Cassini experiment. Their key parameters, along with those of the missions already discussed, are summarized in Table 1.
The MORE, which is on board of the BepiColombo mission launched in 2018 and planned to start orbiting Mercury in 2025 ($r_{r}=0.39$ AU), was designed to provide accurate tracking of the Mercury Planetary Orbiter MORE_1 . The mission aims at studying Mercury’s gravitational field to better understand its internal structure, as well as to perform precision tests of general relativity. For the latter, a similar strategy as for the Cassini mission is followed and two carriers are used for the up- and downlinks: X-band (7.2/8.4 GHz) and Ka-band (32/34 GHz) MORE_1 . With much improved ranging accuracy, it is projected that $|\gamma-1|\lesssim 2\times 10^{-6}$ will be achieved. Plugging these values in eq. (25) we get (assuming $r_{0}=2R_{\odot}$)
[…]
As in the case of the Cassini data, the lower frequency delivers the best results.
$f\,\left[{\rm GHz}\right]$ $r_{e}\,\left[{\rm AU}\right]$ $r_{r}\,\left[{\rm AU}\right]$ $r_{0}\,\left[R_{\odot}\right]$ $|\gamma-1|$
Mariner 6 $2.2$ $1$ $1.7$ $3.6$ $3\times 10^{-2}$bigstrut[t]
Mariner 7 $2.2$ $1$ $1.6$ $5.9$ $3\times 10^{-2}$bigstrut[t]
Viking 1,2 $2.2$ $1$ $1.52$ $1.5$ $2\times 10^{-3}$bigstrut[t]
Cassini 7.2-8.4/32-34 $1$ $8.43$ 1.6 $2.3\times 10^{-5}$bigstrut[t]
MORE 7.2-8.4/32-34 $\approx 1$ $0.39$ $2$ $2\times 10^{-6}$bigstrut[t]
GTDM $2.8\times 10^{5}$ $\approx 1$ $\approx 1$ $2$ $2\times 10^{-8}$bigstrut[t]
Table 1: Key parameters from selected future space-based missions aiming at improving the accuracy in measurements of Eddington’s $\gamma$ parameter (via the Shapiro effect). The data on the Mariner, Viking and Cassini missions are displayed for completeness. For MORE and GTDM we assume $r_{0}=2R_{\odot}$ when the respective transponder spacecraft are closest to superior conjunction; furthermore, the respective emitter spacecraft will be at the L1 Lagrange point $\approx 1.5\times 10^{6}$ km away from Earth in the direction of the Sun, thus $r_{e}\approx 1$ AU.
The GTDM GTDM will use interferometric techniques with $1064$ nm (282 THz) lasers to measure time delays. In fact, lasers are the current golden standard for ranging due to improved accuray and relative compactness of emitter and receiver systems Dirkx . The GTDM will send signals between two drag-free spacecraft – one located near the L1 Lagrange point between Earth and the Sun and the other in orbit at a heliocentric distance of $\approx 1$ AU – to measure time delays to within $0.4$ ps Turyshev . The expected accuracy will be enough to determine $|\gamma-1|$ at the level of $2\times 10^{-8}$, which in turn could be used to place the upper bound (assuming $r_{0}=2R_{\odot}$)
$m_{\gamma}\lesssim 3.0\times 10^{-4}\,{\rm eV/c^{2}}\,.$ (28)
This is significantly worse than the limits extracted from the data of much older missions or the projected bounds from MORE, despite of the higher accuracy in determining $\gamma$ from the GTDM. This is a direct consequence of the use of a high-frequency optical laser, which suppresses the sensitivity by a relative factor of $\approx 10^{5}$.
Figure 2: Bounds on the photon mass versus reported accuracy in the determination of Eddington’s $\gamma$ parameter (expressed as its deviation from unity). In blue are the bounds extracted from data of already completed missions; in red are projected sensitivities based on the design parameters of future time-delay measurements (cf. Table 1).
IV Concluding remarks
In this paper we analysed the gravitational time delay when the electromagnetic signal is composed of very light, but nonetheless massive, photons. Using heliocentric isotropic coordinates we worked out the equations of motion and, following the actual geodesic path, we obtained the two-way (proper) time delay to first post-Newtonian order, cf. eq. (II.2). By comparing this result with the corresponding PPN expression containing Eddington’s $\gamma$ parameter, eq. (18), we were able to extract upper bounds on the photon mass using the actual or projected limits on $|\gamma-1|$ from past and future time-delay measurements. The limits are summarized in Fig. 2.
From the standard PPN expression (18), we see that the time delay due to gravity increases as the light ray gets closer to the Sun (i.e., as $r_{0}$ decreases). Nonetheless, the sensitivity (21) does not benefit from this, as is evident from Fig. 1, where we see the emblematic logarithmic slope of the Shapiro effect acting to worsen the bound for smaller perihelion distances. This somewhat counter intuitive behavior is due to the fact that finite-mass effects, encoded in $\alpha=m_{\gamma}c^{2}/E$, modify only the unobservable Newtonian term and the almost constant post-Newtonian contribution, but not the logarithmic term in eq. (II.2); the correction to go from coordinate to proper time does not matter here. Since the logarithmic term is responsible for the actual time dependence of the Shapiro effect, getting closer to the solar limb means that the $\alpha$-dependent correction is suppressed, thereby worsening the limits on the photon mass.
The upper bounds stemming from data of the Cassini mission represent our best result, reaching the level of $m_{\gamma}\lesssim 3\times 10^{-7}\,{\rm eV/c^{2}}$, cf. eq. (26a). If the technical requirements and mission objectives of the MORE are fulfilled the limit may improve by a factor of $\sim 4$, cf. eq. (27a). Quite generally, our limits improve with decreasing frequency, an effect also reported elsewhere Accioly1 ; Accioly2 ; LSW . This is expected, since smaller masses are related to larger de Broglie wavelengths, meaning that low-frequency phenomena are better probes to finite-mass effects.
@@ IV Concluding remarks
As discussed at the end of Sec. II.1, bounds of similar magnitude were reported in Refs. Accioly1 ; Accioly2 by using measurements of the deflection angle of $\sim 2-43$ GHz radio waves. Though the frequencies used were roughly the same, the accuracy in the determination of $\gamma$ were lower ($\sim 10^{-3}-10^{-4}$ Fomalont ; Lebach ) than what we have used. Nevertheless, their best upper bound is at the same level as ours from the Cassini data, cf. eq. (26a). The issue is, once again, which terms of the standard massless expression are modified by the photon mass. In the case of the deflection angle we have a deviation from general relativity $\sim\delta_{\rm E}\alpha^{2}$, cf. eq. (9). Therefore, any experimental bound will be directly proportional to the photon mass (squared), improving its sensitivity with even modest accuracies in $\gamma$. The situation with the Shapiro effect is unfortunately different, since the main phenomenological contribution, the logarithmic term in eq. (II.2), is not affected by finite-mass corrections. Consequently, the derived limits are not as strong as those from less precise measurements of the bending of radio waves by the Sun (see also the discussion below eq. (19)).
Future space-based missions will likely employ laser signals as a standard, which makes sense since optical frequencies are less impacted than radio by plasma effects in the solar corona, as these scale with $1/f^{2}$ Cassini_1 . Unfortunately, the use of optical lasers is less than ideal in the search for better limits on the photon mass, as is clear from the much better bounds from Cassini and MORE in comparison to the much more precise GTDM, cf. Fig. 2. From eq. (21) we see that radio links with frequencies at the MHz range would be sufficient to reach sensitivities around $m_{\gamma}\lesssim 10^{-13}\,{\rm eV/c^{2}}$ if $|\gamma-1|$ is determined at $10^{-8}$ accuracy, thereby approaching the realm of current bounds based on Earth-scale phenomena PDG ; Kroll ; Malta . In general, the lower the frequency, the better, and if technological limitations are overcome, time-delay measurements at sub-MHz-frequencies in space could potentially surpass other bounds obtained at solar-system scales Ryutov .
@@ Acknowledgements.
We thank J.A. Helayël and Gustavo P. de Brito for reading the manuscript. P.C.M. is indebted to Marina and Karoline Selbach for insightful discussions. This work was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. C.A.D.Z. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under the grant no. 310703/2021-2, and by Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (Faperj) under Grant no. E-26/201.447/2021 (Programa Jovem Cientista do Nosso Estado).

### v2

@@ III Bound on the photon mass via Cassini data
The best measurement of the Shapiro effect was performed using the Cassini spacecraft on its way to Saturn between 6 June and 7 July 2002, when the spacecraft was at heliocentric distances $r_{r}\approx 7.42$ AU [^5] In this period Cassini’s heliocentric distance varied very slightly, linearly going from $r_{r}=7.38$ AU to $r_{r}=7.46$ AU. Here we shall take the constant value $r_{r}=7.42$ AU as a very good approximation. These data were obtained using the JPL Horizons on-line solar system data and ephemeris computation service: https://ssd.jpl.nasa.gov/horizons/app.html.. During its superior conjunction on 21 June 2002 (13:14h UTC), the Cassimi spacecraft was at a geocentric distance of $8.4$ AU and the electromagnetic signal reached its closest approach to the Sun at $r_{\rm 0,min}=1.6\,R_{\odot}$. The radio link was setup so that essentially two carrier signals were sent to and from the spacecraft. The uplink had $f_{u}=7.2$ GHz in the X-band and $f_{u}=34$ GHz in the Ka-band, whereas the downlink had $f_{d}=8.4$ GHz in the X-band and $f_{d}=32$ GHz in the Ka-band. The possibility of using different frequencies allowed the almost complete elimination of coronal effects. Doppler-tracking of the signal was used to constrain $\gamma$ to unprecedented – and so far unsurpassed – accuracy: $|\gamma-1|=2.3\times 10^{-5}$ Cassini_1 ; Cassini_2 .
Instead of trying to directly measure the time delays, the strategy was to observe the difference in frequency of the transmitted (T) and received (R) radar signals. More concretely, the main observable was the fractional frequency shift (FFS) Cassini_1 (cf. eqs. (II.2) and (19))
$y=\frac{f_{\rm R}(\tau)-f_{\rm T}}{f_{\rm T}}=-\frac{d\left(\Delta\tau_{\rm delay}\right)}{d\tau}\,.$ (20)
The data-taking period was roughly a month long and was centered around superior conjunction, so that, as time passes, the main time-dependent parameter is the point of closest approach of the radio signal to the Sun, $r_{0}=r_{0}(\tau)$. Close to Cassini’s superior conjunction, when the spacecraft, the Sun and Earth (in this order) were almost perfectly aligned, the variation of $r_{0}$ was not very different from Earth’s orbital velocity of $30$ km/s. In fact, from Fig. 1 of Ref. Cassini_2 we have
$r_{0}(\tau)=\sqrt{v^{2}\tau^{2}+r^{2}_{0,{\rm min}}}\,$ (21)
with $r_{0,{\rm min}}=1.6\,R_{\odot}$, matching Fig. 2 of Ref. Cassini_imp_param for $v=24.7$ km/s, cf. Fig. 1. Now, $dr_{0}/d\tau=\mathcal{V}(r_{0})$ can be written as
$\mathcal{V}(r_{0})=v\sqrt{1-\left(\frac{r_{0,{\rm min}}}{r_{0}}\right)^{2}}\,,$ (22)
so that
$y=-\mathcal{V}(r_{0})\frac{d\left(\Delta\tau_{\rm delay}\right)}{dr_{0}}\,.$ (23)
Note that $dr_{0}/d\tau<0$ before superior conjunction and $dr_{0}/d\tau>0$ after it, so the sign of $v$ must be changed accordingly.
Our main result, eq. (II.2), can be split in two pieces: one with $\alpha$-independent terms equivalent to the standard predictions of general relativity and one $\alpha$-dependent. Inserting it into eq. (23) the theoretical FFS becomes $y_{\rm th}(\alpha^{2})=y_{0}+\alpha^{2}y_{\alpha}$. The standard contribution is
$y_{0}=\frac{8GM_{\odot}\mathcal{V}(r_{0})}{c^{3}r_{0}}\left[1+W(r_{0},r_{e})+W(r_{0},r_{r})\right]\,,$ (24)
where $8GM_{\odot}\mathcal{V}(r_{0})/c^{3}R_{\odot}\lesssim 10^{-9}$. The leading term stems from the logarithm and the sub-dominant ones from the third line in eq. (II.2). These are controlled by the function
$W(r_{0},r)=\frac{r_{0}r}{2(r+r_{0})^{3/2}\sqrt{r-r_{0}}}\,,$ (25)
shown in Fig. 2. Since $W(r_{0},r)\lesssim\mathcal{O}(10^{-2})$, we are allowed to keep only the leading contribution in $y_{0}$. Finally, the novel, mass-dependent correction to the FFS is
$y_{\alpha}=\frac{8GM_{\odot}\mathcal{V}(r_{0})}{c^{3}r_{0}}\left[W(r_{0},r_{e})+W(r_{0},r_{r})\right]\,.$ (26)
To obtain these expressions we neglected quadratic terms in $r_{0}$ in eq. (II.2), since $r_{0}^{2}\ll r_{e,r}^{2}$.
Figure 1: Impact parameter as a function of time, cf. eq. (21) with $r_{0,{\rm min}}=1.6\,R_{\odot}$ and $v=24.7$ km/s. The curve matches the one in Fig. 2 of Ref. Cassini_imp_param .
Figure 2: The function $W(r_{0},r)$, cf. eq. (25), controls the magnitude of the mass-dependent correction to the FFS. In our current scenario, where $r_{0}\ll r_{e,r}$, we have an approximately linear behavior given by $W(r_{0},r)\approx 2.3\times 10^{-3}(r_{0}/r)$ with $r_{0}$ expressed in units of $R_{\odot}$ and $r$ in AU.
The analysis of the time-delay data performed in Ref. Cassini_2 was based on the accurate characterization of Cassini’s orbital dynamics. This was done using the Orbit Determination Program from NASA’s Jet Propulsion Laboratory with which a fit of the data with up to 12 free parameters, including the initial state vector of the Cassini spacecraft, non-gravitational accelerations, thermo-optical properties of the antenna and Eddington’s $\gamma$ parameter, was performed. The orbital parameters were determined with uncertainties in position $\lesssim 50$ km for geocentric distances of $\approx 8.4$ AU. Furthermore, the uncertainty in the FFS, given by the Allan deviation, was found to be $\sigma_{y}\approx 2\times 10^{-14}$, practically constant and a factor $\approx 10^{4}$ smaller than the measured and predicted FFS Cassini_1 ; Cassini_2 ; Cassini_imp_param ; Cassini_error_budget .
Testing new theories means that their free coefficients must be constrained using the available experimental data – in our case by repeating the data analysis leading to the constraint on $\gamma$. This is particularly true for alternative metric theories of gravity (see e.g. Ref. Hess ), in which the metric tensor and all its by-products, such as the geodesic equations and the resulting orbits, are modified. Massive electrodynamics, however, does not belong to this category: a massive photon, being exquisitely light, does not sufficiently deform spacetime – it acts as a test particle moving in the gravitational field of the Sun and the Schwarzschild metric remains independent of $m_{\gamma}$. Consequently, the orbital dynamics of the Cassini spacecraft (weighing $\approx 5.2$ tons) is unaffected by a tiny, but finite photon mass and the available data may be used without a re-evaluation of the fits.
The data comprise $N=1279$ Doppler (frequency) residuals in the form $\delta f=f_{\rm R}^{\rm obs}-f_{\rm R}^{\rm th}$, where the superscripts “th” and “obs” stand for theoretical and observed, respectively. These are normally distributed about zero with $|\delta f|\lesssim 5\times 10^{-4}$ Hz relative to a carrier frequency $f_{\rm T}=8.4$ GHz. The data are shown in Fig. 3. From eq. (20) we obtain the observed FFS as
$y_{\rm obs}=y_{0}+\frac{\delta f}{f_{\rm T}}\,$ (27)
that may now be used to evaluate the chi-square
$\chi^{2}(\alpha^{2})=\sum_{i=1}^{N}\frac{\left[y_{\rm obs}^{i}-y_{\rm th}^{i}(\alpha^{2})\right]^{2}}{\sigma_{y}^{2}}\,$ (28)
with $\sigma_{y}=2\times 10^{-14}$. The chi-square does not have a minimum for any finite value of $\alpha^{2}$: for any $\{\alpha^{2}_{1},\alpha^{2}_{2}\}$ satisfying $\alpha^{2}_{1}>\alpha^{2}_{2}$ we have $\chi^{2}(\alpha^{2}_{1})>\chi^{2}(\alpha^{2}_{2})$ with a limiting value $\chi_{\rm lim}^{2}/N\approx 0.687$. This indicates that the true minimum is at $\alpha^{2}=0$, that is, the data are compatible with massless electrodynamics, as expected.
We are thus only able to set upper bounds on $\alpha^{2}$. This may be done by searching for $\alpha^{2}=\alpha^{2}_{\star}$ for which $\chi^{2}(\alpha^{2}_{\star})-\chi_{\rm lim}^{2}=3.84$, corresponding to a $95\%$ confidence level PDG . We find $\alpha^{2}_{\star}=1.9\times 10^{-4}$ which translates to (cf. eq. (5))
$m_{\gamma}<4.9\times 10^{-7}\,{\rm eV/c^{2}}\,,$ (29)
where we used $E=hf_{\rm T}=0.035$ meV.
Figure 3: Frequency residuals relative to a carrier frequency $f_{\rm T}=8.4$ GHz and respective impact parameters, courtesy of P. Tortora (U. Bologna). The arrows indicate the time sequence of the measurements from 6 June to 7 July 2002.
IV Concluding remarks
@@
As mentioned at the end of Sec. II.1, bounds of similar magnitude were reported in Refs. Accioly1 ; Accioly2 by using measurements of the deflection angle of $2-43$ GHz radio waves. Though the frequencies used were roughly the same, the experimental uncertainties were larger, also reflecting on the inferior limits on $|\gamma-1|$ Fomalont ; Lebach . Nevertheless, their best upper bound is at the same level as ours, cf. eq. (29). This is explained by looking at which terms of the standard massless expression are modified by the photon mass: in the case of the deflection angle the deviation from general relativity is $\sim\delta_{\rm E}\alpha^{2}$, cf. eq. (10). The situation with the Shapiro effect is unfortunately different, since the main phenomenological contribution, the logarithmic term in eq. (II.2), is not directly affected by finite-mass corrections. Therefore, the very precise measurements performed in the Cassini experiment were crucial to partly compensate for the reduced sensitivity to the photon mass.
Besides the obvious advantage of more precise measurements, one of the most important factors in improving the upper bound on the photon mass is the frequency of the signal. Assuming that $\delta f/f_{\rm T}\approx\sigma_{y}$ remain the same as in the Cassini experiment, from eq. (5) we see that our limits improve with decreasing frequency, an effect also reported elsewhere Accioly1 ; Accioly2 ; LSW . This is in fact expected, since smaller masses are related to larger de Broglie wavelengths, meaning that low-frequency phenomena are better probes to effects related to a very small, but nonetheless finite photon mass.
Looking towards the future, there are a few promising proposals, either already underway or planned, aiming at improving the current limits on various PPN parameters Turyshev . One example is the Mercury Orbiter Radioscience Experiment (MORE) MORE ; MORE_1 on board of the BepiColombo mission launched in 2018 and planned to start orbiting Mercury in 2026 ($r_{r}=0.39$ AU). The objective is to better understand Mercury’s internal structure, as well as to perform precision tests of general relativity. For the latter, a similar strategy to remove coronal effects as in the Cassini mission is followed and two carriers are used. With much improved ranging accuracy and an Allan deviation $\sigma_{y}\approx 10^{-16}$, it is projected that $|\gamma-1|\lesssim 2\times 10^{-6}$ will be achieved. To roughly estimate the impact of these improvement on the upper limit of the photon mass, we may assume that the experimental data still match the predictions from general relativity (to the available precision) and $\delta f/f_{\rm T}\approx\sigma_{y}$ with $f_{\rm T}=8.4$ GHz. This gives us $\alpha^{2}y_{\alpha}\lesssim\sigma_{y}$, meaning that
$m_{\gamma}\sim f_{\rm T}\sqrt{\sigma_{y}}\,,$ (30)
and we may naively expect an improvement of roughly an order of magnitude over eq. (29) just from the projected lower uncertainty.
Future space-based projects, such as the Gravitational Time-Delay Mission Turyshev ; GTDM ; Dirkx , will employ lasers, the current golden standard for ranging due to improved accuray and compactness of emitter and receiver systems Dirkx . This choice is sensible, as optical frequencies are less impacted than radio by coronal plasma effects. Unfortunately, the use of lasers is less than ideal in the search for better limits on the photon mass. From our rough estimate, cf. eq. (30), we see that only substantial improvements in the uncertainties – or conditions which drastically enhance $y_{\alpha}$ (cf. eq. (26)) – will be able to compensate for the $\sim 100$ THz frequencies of optical lasers. Therefore, in order to approach $m_{\gamma}\lesssim 10^{-14}\,{\rm eV/c^{2}}$, the realm of current bounds based on Earth-scale phenomena PDG ; Kroll ; Malta , at realistic error levels, technological limitations must be overcome to allow time-delay measurements at sub-MHz-frequencies in space.
@@ Acknowledgements.
We are grateful to P. Tortora for kindly providing the Cassini time-delay data and patiently clarifying how to use them, and J. Heeck for the support with the statistical analysis. We also thank the anonymous referee for very constructive criticism, M.O. Calvão and L.T. Santana for discussions on the limit of geometric optics, as well as J.A. Helayël and G.P. de Brito for reading the manuscript. P.C.M. and C.A.D.Z. are indebted to Marina and Karoline Selbach, and Lalucha Parizek and Crispim Augusto, respectively, for insightful discussions. This work was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. C.A.D.Z. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) under the grant no. 310703/2021-2, and by Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (Faperj) under Grant no. E-26/201.447/2021 (Programa Jovem Cientista do Nosso Estado).